# DRC: Enhancing Personalized Image Generation via Disentangled Representation Composition

### v1

@@ A.1. Examples of Contrastive Samples
As shown in Figure 6, we display four examples of the contrastive samples. The top two examples correspond to the sticker scenario, while the bottom two illustrate the movie poster scenario.
  • These examples demonstrate that the style-positive images are visually similar to the anchor images. In the sticker scenario, the characters are identical, while in the movie scenario, the images share similar colors and compositional layouts.
  • The semantic-positive images also exhibit strong semantic alignment with the anchor. In the sticker case, the images share similar expressions and movements, whereas in the movie poster case, they feature the same key character from the film.
Figure 6. Four examples of contrastive samples are shown. In each case, the anchor image is displayed in the center. The left two images represent style-positive examples, while the right two images are semantic-positive examples.
@@ A.2. Examples of DRC
As shown in Figure 7, DRC effectively captures the character identity and overall visual style of the history images. At the same time, it accurately reflects the character’s pose and the semantic intent conveyed by the reference image. By integrating both aspects, DRC is able to generate personalized stickers.
Figure 7. Stickers generated by DRC. Dali extracts the style features from the user-interacted history images and the semantic features from the reference image.

### v2

@@ A.1. Examples of Contrastive Samples
In Figure 6, we present several contrastive samples for both sticker and movie poster scenarios, where positive style images share similar style with the anchor but differ in semantics, while positive semantic images share the same semantics but differ in style.
Figure 6. Examples of contrastive samples.
@@ A.2. Examples of DRC
As shown in Figure 7, DRC effectively captures the character identity and overall visual style from the history images. Meanwhile, it accurately reflects the character’s pose and the semantic intent conveyed by the reference image.
Figure 7. Stickers generated by DRC.
A.3. Computational Costs
We present a comparison of training and inference costs between DRC and Pigeon in Table 4. Although DRC introduces an additional disentanglement stage (Stage-1), resulting in a slightly longer total training time compared to Pigeon, it is more memory-efficient during training and achieves significantly faster inference. Overall, we believe this trade-off is favorable, particularly given that inference efficiency is often the key bottleneck in real-world deployment. More importantly, DRC effectively mitigates the guidance collapse issue, as demonstrated in the experimental results in Section 3.
Table 4. Comparison of computational costs.
DRC Pigeon
GPU Memory 55G 60G
Training Stage-1 15.5h -
Stage-2 7.5h 20h
Inference per Sample 7s 16s
A.4. Human Evaluation
[…]